6. A 9-V battery is connected through a switch to
Calculate using physics
Answer #1

Its given that,

Battery Voltage = Vb = 9 volts ; R1 = 15 Ohm and R2 = 23 Ohm ;

L = 0.46 Henry ; V(L) at t = 0.008 seconds

1/Req = 1/R1 + 1/R2 = 1/15 + 1/23

Req = 9.08 Ohm

time constant = L/R = 0.46 / 9.08 = 0.0506

we know that, Voltage across an inductor at time t is given by

V(L) = Vb e-tR/L

V(L) = 9 x e-0.008/0.0506

V(L) = 0.854 Volts